# VizAbility: Enhancing Chart Accessibility with LLM-based Conversational Interaction

### v2

@@ Abstract.
Traditional accessibility methods like alternative text and data tables typically underrepresent data visualization’s full potential. Keyboard-based chart navigation has emerged as a potential solution, yet efficient data exploration remains challenging. We present VizAbility, a novel system that enriches chart content navigation with conversational interaction, enabling users to use natural language for querying visual data trends. VizAbility adapts to the user’s navigation position for improved response accuracy and facilitates verbal command-based chart navigation. Furthermore, it can address queries for contextual information, designed to address the needs of visually impaired users. We designed a large language model (LLM)-based pipeline to address these user queries, leveraging chart data & encoding, user context, and external web knowledge. We conducted both qualitative and quantitative studies to evaluate VizAbility’s multimodal approach. We discuss further opportunities based on the results, including enhanced benchmark testing and integration with current visualization tools.
@@
Figure 1. VizAbility’s overall user interface presents the keyboard navigation of chart content or data tables augmented with the ability to ask natural language questions. Example queries are shown on the right side, including visual & data queries, navigation, and context-seeking quetions.
@@
Figure 1. VizAbility’s overall user interface presents the keyboard navigation of chart content or data tables augmented with the ability to ask natural language questions. Example queries are shown on the right side, including visual & data queries, navigation, and context-seeking quetions.
@@ 1. Introduction
Numerous assistive technologies have been developed to allow BLV users to access visualizations using sensory modalities other than vision (Kim et al., 2021). Tactile visualizations can provide a tangible representation of data while necessitating specialized hardware such as haptic displays  (Paneels and Roberts, 2009) and embossing machines (Engel and Weber, 2017a). On the other hand, sonification can enable users to discern trends and anomalies through sound (Wang et al., 2022), but it is typically limited to single-series data. Traditional methods for adapting web visualizations for screen readers include data tables and alternative text (Kim et al., 2021), while these methods often diminish the inherent advantages of data visualizations. Navigate chart structures with keyboards (Weninger et al., 2015; Elavsky et al., 2023; Thompson et al., 2023; Zong et al., 2022) has emerged as a promising solution. However, a recent comparative study indicates that efficient data exploration, as well as orientation and navigation within complex chart structures, remain significant challenges (Kim et al., 2023a).
@@
This work introduces VizAbility, a novel approach to augment keyboard navigation of chart content with conversational interaction (Figure 1). First, we utilize Olli’s tree view (referenced in (Blanco et al., 2022)) to create a keyboard-navigable text representation of a chart. Next, we enhance this tree view with an LLM-based question-and-answer module for addressing on-demand queries. VizAbility uses the user’s position within the tree view to efficiently respond to both visual and analytical queries, facilitating exploration of data trends and visual patterns. Additionally, it can manage navigation queries, allowing users to control their position through verbal commands. Moreover, VizAbility can handle contextual queries, providing background information about the chart, tailored specifically for the needs of blind and low-vision (BLV) individuals. We based these query types on actual questions asked by BLV people in previous studies (Kim et al., 2020, 2023b). VizAbility’s conversational module concentrates on reasoning with available data and chart encoding information, diverging from existing image-based systems (Masry et al., 2022; Kafle et al., 2018; Kahou et al., 2017). Moreover, unlike the existing similar systems (Murillo-Morales and Miesenberger, 2017; Sharif et al., 2022; Kim et al., 2023b), VizAbility goes beyond basic chart types like bar, line, and pie charts through Vega-lite (Satyanarayan et al., 2016) integration, and address the unique querying needs of BLV individuals with the use of LLM’s advanced language comprehension.
@@
Our LLM-based pipeline first uses few-shot prompting to classify user queries into visual, analytical, contextual, and navigation queries. Once classified, VizAbility employs a query-specific prompting strategy. For analytical and visual queries, we aggregate both the chart’s transformed data and color encoding into one CSV file, which is subsequently fed along with the keyboard-navigable text representation with the user’s location (Blanco et al., 2022) to the LLM via a CSV Agent (csv, [n. d.]). Contextual queries utilize a Web Browser Agent (ser, [n. d.]), whereas navigation queries employ the LLM to discern the starting/ending nodes from a user query and employ a breadth-search algorithm to calculate the shortest path between the nodes. We designed the prompts to minimize hallucinations and address unanswerable queries via structured output formatting. We collaborated with a blind co-design participant in the development of VizAbility, holding two feedback sessions. Their insights, particularly on enhancing interface transparency, were integral to shaping our system design. For example, we implemented strategies to reduce ambiguity in user queries through query refinement and to address potential LLM failures with alternative query suggestions.
[…]
@@
We conducted a usability study with six BLV participants recruited through the National Institute for the Blind. Initially, participants explored VizAbility without guidance and were subsequently introduced to various query types. They also completed the System Usability Scale survey. The results suggest that while participants could learn to use the system, discerning query types without guidance proved challenging. Nonetheless, they acknowledged the merits of the integrated approach and offered suggestions for further improvements and potential applications. For instance, we introduced data tables as an alternative to the tree view and added cold-start query recommendations to assist users in getting started. Combining insights from both quantitative and qualitative evaluations, we identify potential avenues for future work. These include enhancing user-driven customization, developing a more robust benchmarking system, and integrating our solution into existing visualization tools.
@@ 2.1. Accessibility Systems for Data Visualization
Screen readers, utilizing text/speech modalities, stand as the predominant assistive technology, particularly for navigating web content. The go-to accessibility techniques for screen readers encompass alternative text and data tables. Yet, these strategies often reduce data visualizations to brief descriptions or mere numbers, undermining their inherent advantages. An alternative approach involves crafting keyboard-navigable text descriptions derived from the chart’s structure. A select group of data visualization tools and toolkits, such as HighCharts (Highcharts, 2022) and amCharts (amc, [n. d.]), offer some degree of this navigation and customization (Kim et al., 2023a). In recent times, several systems have advanced their navigation capabilities, representing charts as traversable graph structures (Weninger et al., 2015; Godfrey et al., 2018; Zong et al., 2022; Thompson et al., 2023; Elavsky et al., 2023).
@@
In this study, we adopt a different multimodal approach that merges structured chart and table navigation using the keyboard with conversational interaction via verbal commands. Our work builds on the prior work (Kim et al., 2023a) that showcases the respective advantages of data tables—familiarity, structured chart navigation—deeper engagement, and conversational interaction via natural language commands—faster data exploration. Our primary technical advancement centers on employing LLMs to substantially enhance the current chart question-and-answer mechanism for the visually impaired.
@@ 2.2. Question & Answering Systems for Data Visualization
Recent studies have begun to focus on data visualization images (Hoque et al., 2022). For example, FigureQA (Kahou et al., 2017) offers a corpus tailored for yes/no questions, such as “Is Light Gold less than Periwinkle?”. Conversely, DVQA (Kafle et al., 2018) expands its purview to encompass questions about chart structure (“are the bars horizontal?”), data retrieval (“what percent of people prefer A?”), and reasoning (“Is A preferred more than B?”). While both FigureQA and DVQA rely on synthetically generated charts, PlotQA introduces a large-scale dataset of real-world scientific plots. Unlike the templated questions of the aforementioned datasets, ChartQA delivers human-composed questions, enhanced using LLMs (Masry et al., 2022). These models predominantly process pixel images as input. For instance, ChartQA extracts data tables and other image features, feeding them into vision and language task models (Cho et al., 2021). Consequently, their accuracy largely hinges on their image processing capabilities, often leading to suboptimal results. In a different approach, Kim et al.(Kim et al., 2020) unveiled a system that not only answers questions but also provides explanations, operating on Vega-lite(Satyanarayan et al., 2016) instead of images. All the current question-answering systems are limited to basic visualization types like bar, line, and pie charts.
@@
While chart QA systems hint at the potential for enhancing visualization accessibility, they often overlook the specific needs of BLV users. Recent studies have shown that BLV users frame questions differently compared to those with sight (Gurari et al., 2018; Dognin et al., 2020). A limited number of systems directly address the challenge of crafting question-and-answer systems tailored for the blind (Sharif et al., 2022; Murillo-Morales and Miesenberger, 2017). However, these systems do not always offer specialized features for the blind and are constrained in their question-answering capabilities. For instance, VoxLens (Sharif et al., 2022) is limited to charts with single series data, while the system by Murillo-Morales & Miesenberger (Murillo-Morales and Miesenberger, 2017) is restricted to bar charts. Kim et al. (Kim et al., 2023b) have recently curated a set of questions posed by blind individuals through a wizard-of-oz study, laying the groundwork for more refined and targeted question-and-answer systems.
@@
In this paper, we present an enhanced chart question-and-answer system tailored for blind users, leveraging the power of LLMs. Our approach focuses on reasoning, predicated on the availability of chart encoding and underlying data. We utilize Vega-lite as input, thereby accommodating a variety of chart types. The system is adept at handling a wide array of queries. These encompass data and visual queries for delving into complex trends within the chart, as well as contextual queries that necessitate supplementary information about the chart. The basis for these query types is grounded in previous studies ((Kim et al., 2020, 2023b)). Furthermore, our system is designed to support navigation queries to steer keyboard-based chart navigation. We motivated and assessed our system using questions posed by blind individuals, as collected in the study by Kim et al. (Kim et al., 2023b).
@@ 3. VizAbility Design Decisions
In this section, we outline the key design decisions made during the development and evaluation of VizAbility. These decisions were guided by a combination of prior empirical research findings and practical considerations based on user feedback throughout the design process. D: Decisions relevant to our primary contributions D: User interface and usability decisions.
D1: Enable understanding the chart structure
Bridging the perceptual gap between BLV and sighted individuals requires a deep understanding of chart structures. While some blind individuals may not prioritize visual encoding information (Thompson et al., 2023; Lundgard and Satyanarayan, 2021), previous research indicates that navigating charts based on their visual encoding helps BLV users gain a clearer visual understanding. In this work, we utilize Olli (Blanco et al., 2022) to generate a keyboard-navigable text representation of charts. We have further enhanced this tool to include support for choropleth maps.
@@
Furthermore, extracting aggregate measures and discerning perceptual patterns beyond basic value retrievals becomes challenging when navigating data points individually via keyboard input (Kim et al., 2023a). In this work, we adopt a conversational interaction approach that transcends traditional methods such as sonification and tactile perception, which are limited in scalability for modern data visualizations. Leveraging LLMs and user context within keyboard navigation, we address visual and analytic queries that facilitate rapid exploration of nuanced trends and patterns in charts.
@@
Current chart question and answering systems often neglect the distinct types of questions posed by blind versus sighted individuals. Recent research involving blind participants indicates that they frequently ask contextual questions alongside data-related and visual inquiries (Kim et al., 2023b). These questions often seek external information not present in the chart, such as meanings about axes or specific data labels. Providing answers to these inquiries can enhance the self-efficacy and autonomy of blind individuals. In our approach, we utilize an LLM with web search capabilities to address these contextual queries.
@@
Navigating complex chart structures can become less intuitive and more cumbersome (Kim et al., 2023a; Zong et al., 2022), when restricted to keyboard inputs alone. In our work, we aim to mitigate this challenge by facilitating nonlinear investigation across the chart structure via speech commands. In our work, we address this challenge by enabling nonlinear exploration of chart structures through speech commands. This multimodal approach enhances flexibility and efficiency. Furthermore, we aim to assist users in orienting themselves within the chart structure. Understanding one’s position within a digital chart holds equal importance to spatial awareness in physical mobility (Chundury et al., 2021).
@@ 4.1. Input Chart Format
VizAbility operates on the premise that both the visual encoding information and the underlying dataset are accessible. In our work, we employ Vega-Lite specifications (Satyanarayan et al., 2016) as the primary input for our system. Other chart specifications like Observable Plot (obs, [n. d.]) and HighCharts (Highcharts, 2022) can be adapted, provided they expose the underlying data and visual encoding variables, such as axes and legends. These variables are parsed to create a keyboard-navigable tree view, while also providing helpful information about the chart’s appearance for the question-and-answer pipeline, which we detail below.
@@ 4.2. Exploring Chart Encoding using Keyboard
Olli’s tree view displays the chart content in a hierarchical structure, starting with the chart type description at the root—A bar chart. With axes Year and Temperature Anomaly (°C), followed by visual encoding channels such as axes and legends—Legend titled Temporal Polarity. For a nominal scale. With 2 values from negative to positive. Within each encoding channel node, Olli lists data categories or numerical ranges depending on the data type being encoded; e.g., for a color legend, it lists all categories in the legend—1 of 2. Temporal Polarity equals negative. 101 values. Press t to open table. Individual data points reside in these group nodes. All four chart types we used in this work, including line chart, bar chart, scatter plot, and choropleth map, had four levels of information granularity.
A user first needs to enter the tree view to explore the content. Based on its hierarchical structure, users can navigate the different levels of the tree view using up and down arrow keys (barchart $\rightarrow$ legend $\rightarrow$ negative polarity) while using left and right arrow keys to navigate sibling nodes within each level (negative polarity $\rightarrow$ positive polarity). In order to access individual data points, Olli requires users to press t to open up a screen-reader-compatible data table. This table shows a subset of the whole data, only displaying data points within the selected category or numerical range.

### v3

@@ Abstract.
Traditional accessibility methods like alternative text and data tables typically underrepresent data visualization’s full potential. Keyboard-based chart navigation has emerged as a potential solution, yet efficient data exploration remains challenging. We present VizAbility, a novel system that enriches chart content navigation with conversational interaction, enabling users to use natural language for querying visual data trends. VizAbility adapts to the user’s navigation context for improved response accuracy and facilitates verbal command-based chart navigation. Furthermore, it can address queries for contextual information, designed to address the needs of visually impaired users. We designed a large language model (LLM)-based pipeline to address these user queries, leveraging chart data & encoding, user context, and external web knowledge. We conducted both qualitative and quantitative studies to evaluate VizAbility’s multimodal approach. We discuss further opportunities based on the results, including improved benchmark testing, incorporation of vision models, and integration with visualization workflows.
@@
Figure 1. VizAbility’s overall user interface presents the keyboard navigation of chart content or data tables augmented with the ability to ask natural language questions. Example queries are shown on the right side, including visual & data queries, navigation, and context-seeking questions.
[…]
@@ 1. Introduction
Numerous assistive technologies have been developed to allow BLV users to access visualizations using sensory modalities other than vision (Kim et al., 2021). Tactile visualizations can provide a tangible representation of data while necessitating specialized hardware such as haptic displays  (Paneels and Roberts, 2009) and embossing machines (Engel and Weber, 2017a). On the other hand, sonification can enable users to discern trends and anomalies through sound (Wang et al., 2022), but it is typically limited to single-series data. Traditional methods for adapting web visualizations for screen readers include data tables and alternative text (Kim et al., 2021), while these methods often diminish the inherent advantages of data visualizations. Keyboard-based chart navigation (Weninger et al., 2015; Elavsky et al., 2024; Thompson et al., 2023; Zong et al., 2022) have emerged as an alternative solution. However, orienting oneself and navigating within complex chart encoding structures pose challenges for efficient data exploration (Kim et al., 2023a).
@@
This work introduces VizAbility, a novel approach to augment keyboard navigation of chart content with conversational interaction (Figure 1). First, we use Olli’s tree view (referenced in (Blanco et al., 2022)) to create a keyboard-navigable text representation of a chart. Next, we enhance this tree view with an LLM-based question-and-answer module for addressing on-demand queries. These natural language queries allow users to understand the chart without needing to navigate and mentally synthesize different parts of the chart to derive insights. VizAbility uses the user’s position within the tree view to efficiently respond to both visual and analytical queries, facilitating the exploration of data trends and visual patterns. Moreover, VizAbility can handle contextual queries, providing background information about the chart, tailored specifically for the needs of BLV individuals (Kim et al., 2023b). We based these query types on actual questions asked by BLV people in previous studies (Kim et al., 2020, 2023b). Additionally, it can manage navigation queries, allowing users to control their position through verbal commands.
@@
Our LLM-based pipeline first uses few-shot prompting to classify user queries into visual, analytical, contextual, and navigation queries. Once classified, VizAbility employs a query-specific prompting strategy. For analytical and visual queries, we aggregate both the chart’s transformed data and color encoding into one CSV file, which is subsequently fed along with the keyboard-navigable text representation with the user’s location (Blanco et al., 2022) to the LLM via a CSV Agent (csv, [n. d.]). Contextual queries utilize a Web Browser Agent (ser, [n. d.]), whereas navigation queries employ the LLM to discern the starting/ending nodes from a user query and employ a breadth-search algorithm to calculate the shortest path between the nodes. We designed the prompts to minimize hallucinations and address unanswerable queries via structured output formatting. We collaborated with a blind participant in the development of VizAbility, holding a series of feedback sessions.
We carried out quantitative assessments to evaluate the question & answering pipeline. We evaluated response quality using a combined dataset of 979 real BLV user questions derived from previous research (Kim et al., 2023b) and 48 synthetically generated navigation queries. Splitting the dataset, 80% was used for testing and 20% for validation. Our query classification achieved an accuracy of 87.39%. The final response evaluation involved a manual assessment by two researchers, followed by a more scalable LLM-based evaluation using GPT4. Both the human and GPT4 assessments followed a 5-point Likert Scale that assigned a value ranging from “Very Poor” to “Very Good” depending on the response’s coherence to the ground truth. For the human evaluation, $77.36\%$ and $5.14\%$ of the responses were rated as “Very Good’ and “Good” respectively, contributing to an overall rate of $82.50\%$. We computed Kendall’s $\tau$ score ($=0.5526$, $p<0.001$) to assess the consistency between the human and GPT4 assessment methods, observing significant alignment. As a baseline comparison, we used the GPT4-based evaluation to assess responses generated by GPT-4 with vision (GPT-4V) on the same test data. We observed the performance of GPT-4V was poorer, with only 27.96% and 10.10% of responses being designated as “Very Good” and “Good” respectively.
@@
We conducted a preliminary usability study with six BLV participants recruited through the National Institute for the Blind. Initially, participants explored VizAbility without guidance and were subsequently introduced to various query types. They also completed the System Usability Scale survey. The results suggest that while participants could learn to use the system, discerning query types without guidance proved challenging. Nonetheless, they acknowledged the merits of the integrated approach and offered suggestions for further improvements and potential applications. For instance, we introduced data tables as an alternative to the tree view and added cold-start query recommendations to assist users in getting started. Combining insights from both quantitative and qualitative evaluations, we identify potential avenues for future work. These include enhancing user-driven customization, developing a more robust benchmarking system, incorporating vision models, and integrating our solution into existing visualization tools.
Our main contributions lie in the following:
  • Design and development of VizAbility, which incorporates an LLM-based question-and-answer module to enhance keyboard navigation of chart content for screen reader users
  • Development of a benchmark dataset comprising ground truths and chart specifications that augment existing questions posed by blind individuals (Kim et al., 2023b), facilitating further research in this area.
  • Preliminary evaluation that demonstrates the effectiveness of VizAbility in comparison to baseline question-and-answer systems, as well as its usability as assessed by blind participants.
The VizAbility source code and dataset are available at [redacted for anonymous review].
@@ 2.1. Accessibility Systems for Data Visualization
Screen readers, utilizing text/speech modalities, stand as the predominant assistive technology, particularly for navigating web content. The go-to accessibility techniques for screen readers encompass alternative text and data tables. Yet, these strategies often reduce data visualizations to brief descriptions or mere numbers, undermining their inherent advantages. An alternative approach involves crafting keyboard-navigable text descriptions derived from the chart’s structure. A select group of data visualization tools and toolkits, such as HighCharts (Highcharts, 2022) and amCharts (amc, [n. d.]), offer some degree of this navigation and customization (Kim et al., 2023a). In recent times, several systems have advanced their navigation capabilities, representing charts as traversable graph structures (Weninger et al., 2015; Godfrey et al., 2018; Zong et al., 2022; Thompson et al., 2023; Elavsky et al., 2024).
@@
In this study, we adopt an integrated approach that merges structured chart and table navigation using the keyboard with conversational interaction via verbal commands. Our work builds on the prior work (Kim et al., 2023a) that suggests the respective advantages of data tables—familiarity, structured chart navigation—deeper engagement, and conversational interaction via natural language commands—faster data exploration. Our primary technical advancement centers on employing LLMs to enhance the current chart question-and-answer mechanism for the visually impaired.
@@ 2.2. Question & Answering Systems for Data Visualization
Recent studies have begun to focus on data visualization images (Hoque et al., 2022). For example, FigureQA (Kahou et al., 2017) offers a corpus tailored for yes/no questions, such as “Is Light Gold less than Periwinkle?”. Conversely, DVQA (Kafle et al., 2018) expands its purview to encompass questions about chart structure (“are the bars horizontal?”), data retrieval (“what percent of people prefer A?”), and reasoning (“Is A preferred more than B?”). While both FigureQA and DVQA rely on synthetically generated charts, PlotQA introduces a large-scale dataset of real-world scientific plots. Unlike the templated questions of the aforementioned datasets, ChartQA delivers human-composed questions, enhanced using LLMs (Masry et al., 2022). These models predominantly process pixel images as input. For instance, they extract data tables and other image features (Masry et al., 2022; Kantharaj et al., 2022), feeding them into vision and language task models (Cho et al., 2021). Consequently, their accuracy largely hinges on their image processing capabilities, often leading to suboptimal results (e.g., failing to recover data values due to the absence of data labels). In a different approach, Kim et al.(Kim et al., 2020) proposed a system that not only answers questions but also provides explanations, operating on Vega-lite(Satyanarayan et al., 2017) instead of images. All the current question-answering systems are limited to basic visualization types like bar, line, and pie charts.
@@
While chart QA systems hint at the potential for enhancing visualization accessibility, they often overlook the specific needs of BLV users. Recent studies have shown that BLV users frame questions differently compared to those with sight (Gurari et al., 2018; Dognin et al., 2022). A limited number of systems directly address the challenge of crafting question-and-answer systems tailored for the blind (Sharif et al., 2022; Murillo-Morales and Miesenberger, 2017). However, these systems do not always offer specialized features for the blind and are constrained in their question-answering capabilities. For instance, VoxLens (Sharif et al., 2022) is limited to charts with single series data, while the system by Murillo-Morales & Miesenberger (Murillo-Morales and Miesenberger, 2017) is restricted to bar charts. Kim et al. (Kim et al., 2023b) have recently curated a set of questions posed by blind individuals through a wizard-of-oz study, laying the groundwork for more refined and targeted question-and-answer systems.
@@
In this paper, we present an enhanced chart question-and-answer system tailored for blind users, leveraging the power of LLMs. Our approach focuses on reasoning, predicated on the availability of chart encoding and underlying data. We utilize Vega-lite as input, thereby accommodating a variety of chart types. The system handles a wide array of queries, including data and visual queries found in existing question-answer systems, as well as contextual and navigation queries specific to chart accessibility.
@@ 3. VizAbility Design Decisions
In this section, we outline the key design decisions made during the development and evaluation of VizAbility. These decisions were guided by a combination of prior empirical research findings and practical considerations based on user feedback throughout the design and development process. D: Decisions relevant to our primary contributions D: User interface and usability decisions.
D1: Enable understanding the chart encoding structure
Bridging the perceptual gap between BLV and sighted individuals requires a deep understanding of what the chart looks like. While some blind individuals may not prioritize visual encoding information (Thompson et al., 2023; Lundgard and Satyanarayan, 2022), previous research indicates that navigating charts based on their visual encoding helps BLV users gain a clearer visual understanding (Kim et al., 2023a). In this work, we use Olli (Blanco et al., 2022) to generate a keyboard-navigable text representation of charts.
@@
Furthermore, extracting aggregate measures and discerning perceptual patterns beyond basic value retrievals becomes challenging when navigating data points individually via keyboard input (Kim et al., 2023a). This issue exacerbates as the hierarchical text representation becomes deeper and wider with complex chart encodings, resulting in hard mental operations (Green and Petre, 1996). In this work, we adopt a conversational interaction approach that transcends traditional methods such as sonification and tactile perception, which are limited in scalability for modern data visualizations. Leveraging LLMs and user context within keyboard navigation, we address visual and analytic queries that facilitate rapid exploration of nuanced trends and patterns in charts.
@@
Current chart question and answering systems often neglect the distinct types of questions posed by blind versus sighted individuals. Recent research involving blind participants indicates that they frequently ask contextual questions alongside data-related and visual inquiries (Kim et al., 2023b). These questions often seek external information not present in the chart, such as meanings about axes or specific data labels. Providing answers to these inquiries can enhance the self-efficacy and autonomy of blind individuals. In our approach, we use an LLM with web search capabilities to address these contextual queries.
@@
Navigating complex chart structures can become less intuitive and more cumbersome (Kim et al., 2023a; Zong et al., 2022), when restricted to keyboard inputs alone. In our work, we aim to mitigate this challenge by facilitating nonlinear investigation across the chart structure via speech commands. In our work, we address this challenge by enabling nonlinear exploration of chart structures through speech commands. This multimodal approach enhances flexibility and efficiency. Furthermore, we aim to assist users in orienting themselves within the chart structure. Understanding one’s position within a digital chart holds equal importance to spatial awareness in physical mobility (Chundury et al., 2022).
@@ 4.1. Input Chart Format
VizAbility operates on the premise that both the visual encoding information and the underlying dataset are accessible. In our work, we employ Vega-Lite specifications (Satyanarayan et al., 2017) as the primary input for our system. Other chart specifications like Observable Plot (obs, [n. d.]) and HighCharts (Highcharts, 2022) can be adapted, provided they expose the underlying data and visual encoding variables. New adapters will need to be written to work with Olli, which currently supports Vega & Vega-Lite, and Observable Plot (obs, [n. d.]). Alternatively, charts can be translated to the Vega-Lite specifications to be directly used in VizAbility. The symbolic representations underlying the chart are parsed to create a keyboard-navigable tree view and to provide useful information about the chart’s appearance for the question-and-answer pipeline, which is detailed below.
@@ 4.2. Exploring Chart Encoding using Keyboard
Olli’s tree view displays the chart content in a hierarchical structure, starting with the chart type description at the root, followed by visual encoding channels such as axes and legends. Within each encoding channel node, Olli lists data categories or numerical ranges depending on the data type being encoded; e.g., for a color legend, it lists all categories in the legend. Individual data points reside in these group nodes. All four chart types we used in this work, including line chart, bar chart, scatter plot, and choropleth map, had four levels of information granularity.
Based on its hierarchical structure, users can navigate the different levels of the tree view using up and down arrow keys (barchart $\rightarrow$ legend $\rightarrow$ negative polarity) while using left and right arrow keys to navigate sibling nodes within each level (negative polarity $\rightarrow$ positive polarity). In order to access individual data points, Olli requires users to press t to open up a screen-reader-compatible data table. This table shows a subset of the whole data, only displaying data points within the selected category or numerical range.